OpenStudy (anonymous):

For f(x)=-4x+2, find f(-1)

OpenStudy (anonymous):

F(x) and F(-1) Do you see the notation and how it relates?

OpenStudy (anonymous):

\[f(\spadesuit)=-4\times \spadesuit+3\] \[f(-1)=4\times (-1)+2\]

OpenStudy (anonymous):

Give satellite a medal if you haven't already done so @Kasie

OpenStudy (anonymous):

f(♠)=−4×♠+3 f(−1)=4×(−1)+2 = -2

OpenStudy (anonymous):

That's not showing all of the steps, you just showed what satellite put and then put the answer there... That's like going: x^2 + 4x - 2 = -2 -sqrt -6

OpenStudy (anonymous):

Thanks @satellite73